# UniGen: A Unified Framework for Textual Dataset Generation via Large Language Models

### v3

@@ 5 Conclusion
In this paper, we proposed UniGen, a unified dataset generation framework powered by LLMs, which addresses key challenges in diversity, accuracy, and controllability. Its innovative modules and features, ensure high-quality, customizable datasets. The extensive experiments demonstrated the effectiveness of UniGen. Moreover, UniGen can be applied in dynamic and evolving benchmarking as well as data augmentation. Our study has obtained many insightful findings, which have laid the foundation for future research.
@@ Appendix A Impact, Limitation, and Improvement
Our proposed framework, UniGen, not only reduces the costs associated with manually creating data and supports dynamic benchmarking and data augmentation but also significantly impacts the data generation field in several key ways:
  • $\triangleright$
    Alleviating resource scarcity. UniGen effectively addresses the shortage of low-resource datasets. For instance, current datasets were predominantly in English, leaving non-English datasets scarce. Moreover, UniGen can help fill the dataset scarcity in some domains, especially some interdisciplinary fields like AI in psychology [8, 86]. This is significant for both domain development and AI fairness.
  • $\triangleright$
    Enhancing model robustness. The diversity and challenges presented by data generated through UniGen help models improve their ability to handle complex and varied real-world data. This, in turn, enhances the models’ generalization capabilities and reliability, especially in scenarios involving data contamination.
  • $\triangleright$
    Expanding research applications. The methodology used in UniGen can be adapted for other modal data generation frameworks. As models capable of handling different modalities or even multimodal data emerge, the research into data generation for these modalities becomes increasingly relevant and impactful.
@@
    From the perspective of error analysis (subsection 4.5). The error analysis identifies primary areas where UniGen can diminish errors to enhance reliability. To address factuality errors, deploying a robust LLM-based agent [10] enhanced with a broader verification toolkit—comprising an extensive database and web access capabilities—is crucial. Furthermore, question errors frequently stem from LLMs’ misinterpretations of dataset descriptions and objectives, a direct consequence of alignment inefficiencies [87]. Implementing a plug-in module that refines human-written dataset descriptions into formats more comprehensible to LLMs could mitigate this issue.
  • $\triangleright$
@@
    From the perspective of downstream applications (subsection 4.7 and subsection 4.8): A significant oversight in our endeavor to establish a universal dataset generation framework was the insufficient focus on adaptability for specific applications. Concerning dynamic benchmarking protocols such as DyVal [12] and DyVal 2 [13], it is vital to ascertain the specific capabilities that these benchmarks aim to evaluate. For example, while the GSM8K is designed to assess reasoning abilities, the current dataset generation paradigm, which leverages descriptions and few-shot examples, may fail to challenge LLMs adequately. Therefore, orienting the generation process to explicitly target the capabilities under evaluation could truly enhance the dynamism of the dataset. Additionally, our findings indicate limited improvements when applying data augmentation to knowledge-intensive datasets like MMLU [73] and TruthfulQA [72]. A more effective approach could involve identifying novel or out-of-distribution (OOD) data that represents unmastered knowledge for LLMs, thereby significantly enhancing learning outcomes.
  • $\triangleright$
@@
    From the perspective of weak-to-strong alignment [88, 89] & self-alignment [16, 90, 91]: LLM-generated data have been extensively utilized to improve LLMs themselves. For example, Phi-3 [81] is trained using a substantial amount of synthetic data generated by GPT-4. This utilization demonstrates that LLMs can undergo self-evolution through synthetic data. In our study, while we have explored potential alignments in a cross-model mode (e.g., using GPT-4 to enhance weaker models), the strategies for self-alignment or weak-to-strong alignment within the same model are not thoroughly investigated. Future research focusing on how to adapt a dataset generation framework like UniGen for use in data-centric alignment domains will be of considerable importance.
Appendix B Details of Datasets and Models
B.1 Datasets
@@
GSM8K. GSM8K is a dataset designed to test the mathematical problem-solving ability of large language models [19]. It comprises approximately 8,000 math word problems typical of those in grade school. The problems are diverse, covering various topics and difficulties, making it a comprehensive tool for assessing the reasoning capabilities of models in numerical contexts.
@@
TruthfulQA. TruthfulQA is a dataset crafted to evaluate the truthfulness and factual accuracy of answers provided by language models [72]. It consists of questions that models frequently respond to incorrectly or misleadingly. The dataset challenges models on simple factual questions and questions requiring a nuanced understanding of common misconceptions and controversial topics.
@@
MMLU. MMLU is a large-scale dataset designed to test various language understanding tasks [73]. It covers 57 subjects ranging from humanities to natural sciences, providing a broad spectrum of topics. This diversity makes MMLU highly effective for assessing the general knowledge and understanding of language models across varied domains.
@@
HellaSwag. HellaSwag is a dataset that evaluates common sense reasoning and context understanding in language models [74]. It includes scenarios requiring the prediction of the most plausible continuation among several options. The dataset is crafted to be particularly challenging, often including subtle nuances and twists that test the depth of contextual comprehension.
@@
MetaTool. MetaTool is a benchmark designed to evaluate LLMs’ awareness and proficiency in tool usage and selection [9]. In our experiment, we conducted evaluations on two tasks. In our experiments, we specifically focused on single-tool selection.
@@
MultiNLI. The Multi-Genre Natural Language Inference (MultiNLI) is a crowd-sourced dataset of 433k sentence pairs annotated with textual entailment information [92]. It covers a range of genres of spoken and written text and supports a distinctive cross-genre generalization evaluation.
@@
ARC (Challenge). The AI2’s Reasoning Challenge (ARC) dataset is a multiple-choice question-answering dataset, containing questions from science exams from grade 3 to grade 9 [93]. The dataset is split into two partitions: Easy and Challenge, where the latter partition contains the more difficult questions that require reasoning.
@@
BoolQ. BoolQ is a reading comprehension dataset with questions that are unexpectedly challenging [94]. They often query for complex, non-factoid information, and require difficult entailment-like inference to solve.
@@
BBH. BIG-Bench Hard (BBH) is a subset of the BIG-Bench, a diverse evaluation suite for language models [84]. BBH focuses on a suite of 23 challenging tasks from BIG-Bench that were found to be beyond the capabilities of current language models. We select two tasks from BBH: boolean expressionshttps://github.com/suzgunmirac/BIG-Bench-Hard/blob/main/bbh/boolean_expressions.json and causal judgementhttps://github.com/suzgunmirac/BIG-Bench-Hard/blob/main/bbh/causal_judgement.json.
Table 8: The dataset description we used in UniGen.
@@
GSM8K It is a dataset of high-quality linguistically diverse grade school math word problems created by human problem writers. These problems take between 2 and 8 steps to solve, and solutions primarily involve performing a sequence of elementary calculations using basic arithmetic operations ($+$ -- $\times$ $\div$) to reach the final answer. A bright middle school student should be able to solve every problem. It can be used for multi-step mathematical reasoning. Each problem should only have one question and one correct answer.
@@
B.2 Models
Models for Benchmarking. These include ChatGPT [1] and GPT-4 [95] by OpenAI [96], known for their robust conversational abilities; Llama3-70b and Llama3-8b [3] by Meta AI [97], open-source and favored for their versatility across different computational scales; Mistral-7b and Mistral-8x7b [98] by Mistral AI [99], designed for efficiency in language tasks; Claude3 [2] by Anthropic [100], which focuses on safe and ethical AI interactions; and Yi-34b [101] by 01.AI [102], a model fine-tuned using high-quality curated data to ensure helpfulness.
Appendix C Details of Human Evaluation
We conduct human evaluations in two parts: effectiveness of each module in UniGen (subsection 4.3) and error analysis (subsection 4.5). Four undergraduate students and one PhD student with professional English skills carry out these evaluations. Some annotation screenshots of human evaluation are shown in Figure 13 and Figure 14.
Effectiveness of Each Module in UniGen. In subsection 4.3, we conduct the human evaluation of overall quality assessment and enhancement, code-based, and RAG-based validation. Specifically, for code-based evaluation, when a label contradicts the code output, we will manually check whether the code output is correct (in UniGen, we will replace the original label with code output if they contradict). For the RAG-based validation, we also manually whether the correcting action is reasonable and supported by the ground truth.
@@ Appendix C Details of Human Evaluation
Error Analysis. The error analysis (subsection 4.5) is based on a structured human evaluation approach. To ensure the quality of the generated questions, human experts review each question against specific criteria that cover various aspects of data integrity and logical coherence. Below are the detailed aspects that are evaluated:
@@
Appendix D Details of Experiment Setting
Dataset Generation. To maximize the consistency of the experimental results, we set the temperature parameter for both GPT-4 and Claude-3 to 0. The size of the generated dataset used in subsection 4.2 and benchmarking LLMs is shown in Table 9. The batch size of generation (the number of items generated per time) is set to 5.
Table 9: The size of the generated dataset used in subsection 4.2 and benchmarking LLMs.
@@ Appendix D Details of Experiment Setting
Inference Settings.Across all models, we maintained uniform hyperparameter settings. Specifically, the model temperature was set to 0 to enhance productivity, and the top-p was set to 1. For benchmarking purposes with Mixtral-8x7b and Llama3-70b, we utilized the inference API provided by Replicate§§§https://replicate.com/.
Fine-tune Settings.For each dataset, UniGen generates 200 samples powered by GPT-4 and then evaluates the fine-tuned models on the test set of the original dataset. The labels or ground-truth answers of generated data always contain only a few words, lacking a thinking process that may be more important for fine-tuning. To address this, the labels or the ground-truth answers of the generated dataset are refined and extended by GPT-4 itself (e.g., transform the answers into Chain-of-Thoughts format [103]). Then a self-evaluation of GPT-4 will be conducted to ensure the correctness and accuracy of refined answers. Our fine-tuning is all based on the Supervised Fine-Tuning (SFT):
@@
We applied the LoRA [104] technique to fine-tune Llama3-8b and Mistral-7b. The rank of LoRA was set to 8, the learning rate was $e^{-5}$, and we used the Adam optimizer [105] for training. The models were trained over 5 epochs with a batch size of 4, utilizing mixed precision training. The training took place on a server equipped with an A100 GPU with 80GB of VRAM. For the training process, we employed the LLAMA-Factory framework [106].
Appendix E Additional Experiment Results
@@
Influence of temperature. We examined the impact of temperature settings on the diversity of data generated by GPT-4. For this purpose, we selected a few items from the TruthfulQA dataset to use as examples in few-shot learning. We conducted experiments using temperature settings of 0 and 1. Our findings indicate that the Remote-Clique score [76] at a temperature of 0 is 0.683, whereas, at a temperature of 1, it increases to 0.721. This suggests that adjusting the temperature setting can significantly enhance the diversity of the generated data.
Figure 9: Semantic embedding of different datasets. We use OpenAI’s text-embedding-ada-002 [59] to obtain text embedding.
User Constraints. To evaluate the effectiveness of LLMs in UniGen at adhering to user-specified constraints, our assessment is structured into two levels. The first level involves evaluating the model’s performance under single constraints, while the second level examines performance under combined constraints. The single constraints assessed include:
@@
Table 11: The GP-4’s performance on user constraints.
@@
[…]
@@
To assess whether the LLM adheres to user-imposed constraints, we utilize the LLM-as-a-Judge approach [46], a method extensively employed in prior research [44, 70]. The evaluation prompt details are provided in Appendix I. As indicated in Appendix E, GPT-4 demonstrates outstanding performance across both single and combined constraints. It achieves a 100% compliance rate in nine out of ten single constraints, illustrating its robust capability to follow simple and typical user instructions. Although there is a slight performance decline in combined constraints, GPT-4 consistently maintains adherence to user constraints in most scenarios.
@@
Appendix F Potential Negative Societal Impacts

### v5

@@ 4 Conclusion
In this paper, we hava proposed DataGen, a unified dataset generation framework powered by LLMs, which addresses key challenges in diversity, accuracy, and controllability. Its innovative modules and features, ensure high-quality, customizable datasets. The extensive experiments demonstrated the effectiveness of DataGen. Moreover, DataGen can be applied in dynamic and evolving benchmarking as well as data augmentation. We believe that the insightful findings revealed in this study will serve as a foundation for future research on data generation.
@@ Appendix A Impact, Limitation, and Improvement
Our proposed framework, DataGen, not only reduces the costs associated with manually creating data and supports dynamic benchmarking and data augmentation but also significantly impacts the data generation field in several key ways:
  • Alleviating resource scarcity. DataGen effectively addresses the shortage of low-resource datasets. For instance, current datasets were predominantly in English, leaving non-English datasets scarce. Moreover, DataGen can help fill the dataset scarcity in some domains, especially some interdisciplinary fields like AI in psychology [li2024i]. This is significant for both domain development and AI fairness.
  • Enhancing model robustness. The diversity and challenges presented by data generated through DataGen help models improve their ability to handle complex and varied real-world data. This, in turn, enhances the models’ generalization capabilities and reliability, especially in scenarios involving data contamination.
  • Expanding research applications. The methodology used in DataGen can be adapted for other modal data generation frameworks. As models capable of handling different modalities or even multimodal data emerge, the research into data generation for these modalities becomes increasingly relevant and impactful.
@@
  • From the perspective of error analysis (Section 3.5). The error analysis identifies primary areas where DataGen can diminish errors to enhance reliability. To address factuality errors, deploying a robust LLM-based agent [liu2023agentbench] enhanced with a broader verification toolkit—comprising an extensive database and web access capabilities—is crucial. Furthermore, question errors frequently stem from LLMs’ misinterpretations of dataset descriptions and objectives, a direct consequence of alignment inefficiencies [ji2024ai]. Implementing a plug-in module that refines human-written dataset descriptions into formats more comprehensible to LLMs could mitigate this issue.
@@
  • From the perspective of downstream applications (Section 3.7 and Section 3.8): A significant oversight in our endeavor to establish a universal dataset generation framework was the insufficient focus on adaptability for specific applications. Concerning dynamic benchmarking protocols such as DyVal [zhu2024dyval] and DyVal 2 [zhu2024dyval2], it is vital to ascertain the specific capabilities that these benchmarks aim to evaluate. For example, while the GSM8K is designed to assess reasoning abilities, the current dataset generation paradigm, which leverages descriptions and few-shot examples, may fail to challenge LLMs adequately. Therefore, orienting the generation process to explicitly target the capabilities under evaluation could truly enhance the dynamism of the dataset. Additionally, our findings indicate limited improvements when applying data augmentation to knowledge-intensive datasets like MMLU [hendrycks2021measuringmassivemultitasklanguage] and TruthfulQA [lin2022truthfulqa]. A more effective approach could involve identifying novel or out-of-distribution (OOD) data that represents unmastered knowledge for LLMs, thereby significantly enhancing learning outcomes.
@@
  • From the perspective of weak-to-strong alignment [zheng2024weak, burns2023weak] & self-alignment [sun2023principledriven, li2023self, sun2024principle]: LLM-generated data have been extensively utilized to improve LLMs themselves. For example, Phi-3 [abdin2024phi3] is trained using a substantial amount of synthetic data generated by GPT-4. This utilization demonstrates that LLMs can undergo self-evolution through synthetic data. In our study, while we have explored potential alignments in a cross-model mode (e.g., using GPT-4 to enhance weaker models), the strategies for self-alignment or weak-to-strong alignment within the same model are not thoroughly investigated. Future research focusing on how to adapt a dataset generation framework like DataGen for use in data-centric alignment domains will be of considerable importance.
Appendix B Related Work
Benchmarking and Evaluating LLMs.
Owing to the remarkable capabilities of LLMs, benchmarking these models is essential for a deeper understanding of both general and specialized domains [chang2023survey]. The evaluation of LLMs encompasses a wide range of fields, initiating with core NLP tasks such as sentiment analysis [lopezlira2023chatgpt, zhang2023sentiment], text classification [yang2023large, zhang2023generationdriven], and natural language inference [mckenna2023sources]. A holistic evaluation framework, the HELM benchmark, has been proposed by liang2023holistic, laying the groundwork for comprehensive assessments. Additionally, the application of LLMs spans diverse sectors [gu2023xiezhi], including computational social science [ziems2023large], legal analytics [nay2023large, guha2023legalbench, fei2023lawbench], and psychological studies [fmc2023, li2024i]. Furthermore, several benchmarks have been designed to scrutinize trustworthiness dimensions such as safety and privacy in LLMs [sun2024trustllm, huang2023trustgpt, wang2024decodingtrust, gao2024honestllm, huang2024social, li2025preference, zhou2024labsafety, huang2025trustworthiness]. Static benchmarks are susceptible to data contamination, wherein developers might incorporate benchmark datasets into the training data to artificially enhance performance. To mitigate this, flexible protocols for dynamic evaluation have been advanced, exemplified by the recent initiatives DyVal [zhu2024dyval] and DyVal 2 [zhu2024dyval2]. Additionally, fan2024nphardeval introduced NPHardEval, featuring monthly updated datasets. The S3Eval framework, a scalable evaluation suite for LLMs, was conceptualized by [lei2024s3eval]. bao2024autobench introduce a framework to automatically evaluate VLLMs by themselves. Moreover, some benchmarks adopt methodologies where LLMs function as evaluators (e.g., LLM-as-a-judge) [liu2023alignbench, chen2024mllmasajudge, zheng2023judging, ye2024justice], with AlignBench proposing a multi-dimensional assessment using this approach [liu2023alignbench].
Synthetic Data by LLMs.
LLMs have demonstrated an impressive capacity for data generation, leading to their application in creating synthetic datasets for pretraining and finetuning, replacing the labor-intensive processes of manual data scraping and selection [liu2024best]. Distinct from earlier methods that focus on traditional language models [schick2021generating], LLMs offer enhanced prospects for producing high-quality synthetic data across a wide spectrum of applications, such as multilingual QA [riabi-etal-2021-synthetic], chatbot conversation [zhao2023inthe], instruction tuning [xu2024magpie], and data diversity augmentation [dai2023auggpt, chung2023increasing, chen2024interleaved].
The concept of synthetic benchmarks takes a step further by demanding that the LLM-generated data be diverse accurate and systematically challenging. For instance, wang2024benchmark devised a framework that enhances the evolution of benchmarks by applying six reframing techniques on existing datasets. wei2024longform employed GPT-4 to create LongFact, comprising extensive QA pairs that serve as a benchmark for evaluating long-form factual content. Moreover, synthetic benchmarks have also been constructed in evaluating LLM emergent capabilities such as trustworthiness [sun2024trustllm, huang2025breaking], tool usage [huang2024metatool, qin2023toolllm] and persona-based conversation [jandaghi2023faithful]. Our research advances synthetic benchmark generation by developing a paradigm that integrates multiple plug-and-play modules into LLM dataset creation, leveraging emergent capabilities by various prompting methods (e.g., self-evaluation [ji2023towards]) to produce data items with high-quality. Recently, in response to concerns about the quality of synthetic datasets, dekoninck2024understanding conducted comprehensive experiments to evaluate the diversity and fidelity of synthetic data produced by LLMs, while dekoninck2024controlled introduced a new inference framework, model arithmetic, to control the content generated by LLMs.
Appendix C Details of Datasets and Models
C.1 Datasets
@@
GSM8K. GSM8K is a dataset designed to test the mathematical problem-solving ability of large language models [cobbe2021gsm8k]. It comprises approximately 8,000 math word problems typical of those in grade school. The problems are diverse, covering various topics and difficulties, making it a comprehensive tool for assessing the reasoning capabilities of models in numerical contexts.
@@
TruthfulQA. TruthfulQA is a dataset crafted to evaluate the truthfulness and factual accuracy of answers provided by language models [lin2022truthfulqa]. It consists of questions that models frequently respond to incorrectly or misleadingly. The dataset challenges models on simple factual questions and questions requiring a nuanced understanding of common misconceptions and controversial topics.
@@
MMLU. MMLU is a large-scale dataset designed to test various language understanding tasks [hendrycks2021measuringmassivemultitasklanguage]. It covers 57 subjects ranging from humanities to natural sciences, providing a broad spectrum of topics. This diversity makes MMLU highly effective for assessing the general knowledge and understanding of language models across varied domains.
@@
HellaSwag. HellaSwag is a dataset that evaluates common sense reasoning and context understanding in language models [zellers2019hellaswag]. It includes scenarios requiring the prediction of the most plausible continuation among several options. The dataset is crafted to be particularly challenging, often including subtle nuances and twists that test the depth of contextual comprehension.
@@
MetaTool. MetaTool is a benchmark designed to evaluate LLMs’ awareness and proficiency in tool usage and selection [huang2024metatool]. In our experiment, we conducted evaluations on two tasks. In our experiments, we specifically focused on single-tool selection.
@@
MultiNLI. The Multi-Genre Natural Language Inference (MultiNLI) is a crowd-sourced dataset of 433k sentence pairs annotated with textual entailment information [williams2018broadcoverage]. It covers a range of genres of spoken and written text and supports a distinctive cross-genre generalization evaluation.
@@
ARC (Challenge). The AI2’s Reasoning Challenge (ARC) dataset is a multiple-choice question-answering dataset, containing questions from science exams from grade 3 to grade 9 [Clark2018ThinkYH]. The dataset is split into two partitions: Easy and Challenge, where the latter partition contains the more difficult questions that require reasoning.
@@
BoolQ. BoolQ is a reading comprehension dataset with questions that are unexpectedly challenging [clark2019boolq]. They often query for complex, non-factoid information, and require difficult entailment-like inference to solve.
@@
BBH. BIG-Bench Hard (BBH) is a subset of the BIG-Bench, a diverse evaluation suite for language models [suzgun2022challenging]. BBH focuses on a suite of 23 challenging tasks from BIG-Bench that were found to be beyond the capabilities of current language models. We select two tasks from BBH: boolean expressions and causal judgement.
Table 9: The dataset description we used in DataGen.
@@
GSM8K It is a dataset of high-quality linguistically diverse grade school math word problems created by human problem writers. These problems take between 2 and 8 steps to solve, and solutions primarily involve performing a sequence of elementary calculations using basic arithmetic operations (++ - $\times$ $\div$) to reach the final answer. A bright middle school student should be able to solve every problem. It can be used for multi-step mathematical reasoning. Each problem should only have one question and one correct answer.
@@
C.2 Models
Models for Benchmarking. These include ChatGPT [ChatGPT] and GPT-4 [GPT-4] by OpenAI [OpenAI], known for their robust conversational abilities; Llama3-70b and Llama3-8b [LLAMA3] by Meta AI [MetaAILab], open-source and favored for their versatility across different computational scales; Mistral-7b and Mistral-8x7b [jiang2024mixtral] by Mistral AI [MistralAI], designed for efficiency in language tasks; Claude3 [Claude] by Anthropic [Anthropic], which focuses on safe and ethical AI interactions; and Yi-34b [ai2024yi] by 01.AI [01AI], a model fine-tuned using high-quality curated data to ensure helpfulness.
Appendix D Details of Human Evaluation
We conduct human evaluations in two parts: effectiveness of each module in DataGen (Section 3.3) and error analysis (Section 3.5). Four undergraduate students and one PhD student with professional English skills carry out these evaluations. Some annotation screenshots of human evaluation are shown in Figure 13 and Figure 14.
Effectiveness of Each Module in DataGen. In Section 3.3, we conduct the human ablation evaluation of overall quality assessment and enhancement, code-based, and RAG-based validation. Specifically, for code-based evaluation, when a label contradicts the code output, we will manually check whether the code output is correct (in DataGen, we will replace the original label with code output if they contradict). For the RAG-based validation, we also manually whether the correcting action is reasonable and supported by the ground truth.
@@ Appendix D Details of Human Evaluation
Error Analysis. The error analysis (Section 3.5) is based on a structured human evaluation approach. To ensure the quality of the generated questions, human experts review each question against specific criteria that cover various aspects of data integrity and logical coherence. Below are the detailed aspects that are evaluated:
@@
Appendix E Details of Experiment Setting
Dataset Generation. To maximize the consistency of the experimental results, we set the temperature parameter for both GPT-4 and Claude-3 to 0. The size of the generated dataset used in Section 3.2 and benchmarking LLMs is shown in Table 10. The batch size of generation (the number of items generated per time) is set to 5.
Table 10: The size of the generated dataset used in Section 3.2 and benchmarking LLMs.
@@ Appendix E Details of Experiment Setting
Inference Settings. We maintained uniform hyperparameter settings across all models. Specifically, the model temperature was set to 0 to enhance productivity, and the top-p was set to 1. For benchmarking purposes with Mixtral-8x7b and Llama3-70b, we utilized the inference API provided by Replicatehttps://replicate.com/.
Fine-tune Settings. For each dataset, DataGen generates 200 samples powered by GPT-4 and then evaluates the fine-tuned models on the test set of the original dataset. The labels or ground-truth answers of generated data always contain only a few words, lacking a thinking process that may be more important for fine-tuning. To address this, the labels or the ground-truth answers of the generated dataset are refined and extended by GPT-4 itself (e.g., transform the answers into Chain-of-Thoughts format [wei2023chainofthought]). Then a self-evaluation of GPT-4 will be conducted to ensure the correctness and accuracy of refined answers. Our fine-tuning is all based on the Supervised Fine-Tuning (SFT):
@@
We applied the LoRA [hu2021lora] technique to fine-tune Llama3-8b and Mistral-7b. The rank of LoRA was set to 8, the learning rate was $e^{-5}$, and we used the Adam optimizer [kingma2017adam] for training. The models were trained over 5 epochs with a batch size of 4, utilizing mixed precision training. The training took place on a server equipped with an A100 GPU with 80GB of VRAM. For the training process, we employed the LLAMA-Factory framework [zheng2024llamafactory].
Appendix F Additional Experiment Results
@@
User Constraints. To evaluate the effectiveness of LLMs in DataGen at adhering to user-specified constraints, our assessment is structured into two levels. The first level involves evaluating the model’s performance under single constraints, while the second level examines performance under combined constraints. The single constraints assessed include:
@@
Table 12: The GPT-4’s performance on user constraints.
@@
Single Constraint ($\uparrow$), Combined Constraint ($\downarrow$)
Constraint 1 Constraint 2 Constraint 3 Constraint 4 Constraint 5
@@
To assess whether the LLM adheres to user-imposed constraints, we utilize the LLM-as-a-Judge approach [zheng2023judging], a method extensively employed in prior research [liu2023alignbench, gao2024best]. The evaluation prompt details are provided in Appendix J. As indicated in Table 12, GPT-4 demonstrates outstanding performance across both single and combined constraints. It achieves a 100% compliance rate in nine out of ten single constraints, illustrating its robust capability to follow simple and typical user instructions. Although there is a slight performance decline in combined constraints, GPT-4 consistently maintains adherence to user constraints in most scenarios.
@@
Diversity. For more features of generated data, we have referred to the study [yu2024large] to guide our incorporation of two quantitative metrics to evaluate dataset diversity: the Average Pairwise Sample Similarity (APS) and the Inter-Sample N-Gram Frequency (INGF). Lower APS values indicate better diversity, whereas higher INGF values signify greater diversity. The result is shown in Table 14.
Table 14: Comparison of Original and Generated APS and INGF values across datasets
Dataset Original APS Generated APS Original INGF Generated INGF
TruthfulQ&A 0.029 0.091 882.181 1603.976
GSM8K 0.053 0.057 3021.619 1296.588
MMLU 0.047 0.050 2185.514 1566.574
HellaSwag 0.076 0.089 2586.710 2193.623
@@
Appendix G Potential Negative Societal Impacts